OpenStudy (anonymous):

Give an example of a function with both a removable and a non-removable discontinuity.

OpenStudy (anonymous):

Could you help? @satellite73

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

removable discontinuity is one of those "factor and cancel" deals, like for example \[\frac{x^2-4}{x-2}\] this is undefined at \(x=2\) but it is also true that if \(x\neq 2\) this is the same as \[\frac{(x+2)(x-2)}{x-2}=x+2\]

OpenStudy (anonymous):

the discontinuity is "removable" because... well because when you factor and cancel you remove it

OpenStudy (anonymous):

and so the non-removable discontinuity would be x+2 ?

OpenStudy (anonymous):

because it isn't removed?

OpenStudy (anonymous):

lol no that is not a discontinuity that is an expression

OpenStudy (anonymous):

so the discontinuity is the entire thing? where parts cancel each other out?

OpenStudy (anonymous):

the discontinuity was at \(2\) and we removed it when we rewrote \[\frac{(x+2)(x-2)}{x-2}=x+2\]

OpenStudy (anonymous):

the discontinuity is at a number \(2\) was the removable discontinuity

OpenStudy (anonymous):

and what does 2 represent?

OpenStudy (anonymous):

Are you there? @satellite73